# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 6      $t\leftarrow t+1$
Output: Coefficients $[(\lambda_{0},\widetilde{\bm{\beta}}(\lambda_{0})),(\lambda_{1},\widetilde{\bm{\beta}}(\lambda_{1})),\dots]$
@@
We now present our approximate quantum LARS algorithm 3 for the pathwise Lasso. Its main idea is quite simple: we improve the search of the joining variable $i_{t+1}^{\rm join}$ and time $\lambda_{t+1}^{\rm join}$ over $i\in\mathcal{I}$ by approximately computing the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and using the approximate quantum minimum-finding subroutine from Chen and de Wolf [CdW23] (Fact 11). More specifically, the joining variable $\widetilde{i}_{t+1}^{\rm join}$ is obtained to be any variable $j\in\mathcal{I}$ that maximises $\Lambda_{j}^{\rm join}(\mathcal{A})$ up to some small relative error $\epsilon$, i.e., at any iteration, Algorithm 3 randomly samples from the set
$\left\{j\in\mathcal{I}:\Lambda_{j}^{\rm join}(\mathcal{A})\geq(1-\epsilon/2)\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\right\}.$ (13)
Since the corresponding joining time $\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A})$ can be smaller than the true joining time $\lambda_{t+1}^{\rm join}=\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}$, we rescale it by the factor $(1-\epsilon/2)^{-1}$ and set $\widetilde{\lambda}_{t+1}^{\rm join}=(1-\epsilon/2)^{-1}\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A})$. This ensures that $\widetilde{\lambda}_{t+1}^{\rm join}\geq\lambda_{t+1}^{\rm join}$ and consequently that the current iteration stops before $\lambda_{t+1}^{\rm join}$, which guarantees that the approximate KKT conditions are satisfied as shown in Theorem 25 below.
We notice that, since the joining variable is randomly sampled from the set in Eq. (13), it might be possible that $\widetilde{\lambda}_{t+2}^{\rm join}>\widetilde{\lambda}_{t+1}^{\rm join}$, i.e., the joining time at some later iteration is greater than the joining time at some earlier iteration. While this is counter to the original LARS algorithm, where the regularisation parameter $\lambda$ always decreases, there is in principle no problem in allowing $\lambda$ to increase within a small interval as long as the approximate KKT conditions are satisfied. Nonetheless, in order to avoid redundant segments in the Lasso path, we set the next iteration’s regularisation parameter $\lambda_{t+1}$ as the minimum between the current $\lambda_{t}$ and $\max\{\lambda_{t+1}^{\rm cross},\widetilde{\lambda}_{t+1}^{\rm join}\}$. The solution path can thus stay “stationary” for a few iterations. Ideally, one could just move all the variables from the set in Eq. (13) into $\mathcal{A}$ at once and the result would be the same. This is reminiscent to the approximate homotopy algorithm of Mairal and Yu [MY12].
@@
Let $\epsilon\in[0,2)$. Consider an approximate LARS algorithm (e.g. Algorithm 3) that returns a solution path $\widetilde{\mathcal{P}}=\{\widetilde{\bm{\beta}}(\lambda)\in\mathbb{R}^{d}:\lambda>0\}$ and wherein, at each iteration $t$, the joining variable $\widetilde{i}_{t+1}^{\rm join}$ is taken from the set,
$\left\{j\in\mathcal{I}:\Lambda_{j}^{\rm join}(\mathcal{A})\geq(1-\epsilon/2)\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\right\},$
where $\Lambda_{j}^{\rm join}(\mathcal{A})\triangleq\frac{\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{j}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$, and the corresponding joining time $\widetilde{\lambda}_{t+1}^{\rm join}$ is
$\widetilde{\lambda}_{t+1}^{\rm join}=(1-\epsilon/2)^{-1}\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A}),$
@@
$\bm{\eta}_{\mathcal{A}}=\frac{1}{\lambda_{t}}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t})),\quad\quad\quad\bm{\mu}=\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y},\quad\quad\quad\bm{\theta}=(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}.$
@@
The proof is by induction on the iteration loop $t$. The case $t=0$ is trivial. Assume then that the computed path through iteration $t-1$ satisfies $\mathsf{KKT}_{\lambda}(\epsilon)$ for all $\lambda\geq\lambda_{t}$. Consider the $t$-th iteration with active set $\mathcal{A}=\{i\in[d]:\widetilde{\beta}_{i}\neq 0\}$ and equicorrelation signs $\bm{\eta}_{\mathcal{A}}=\frac{1}{\lambda_{t}}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda_{t}))$. The induction hypothesis implies that the current Lasso solution $\widetilde{\bm{\beta}}(\lambda_{t})$ satisfies $\mathsf{KKT}_{\lambda_{t}}(\epsilon)$ at $\lambda_{t}$, i.e.,
$\lambda_{t}(1-\epsilon)\leq\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda_{t}))\operatorname{sign}(\widetilde{\beta}_{j}(\lambda_{t}))\leq\lambda_{t}$ $\quad~{}\text{if}~{}\widetilde{\beta}_{j}\neq 0,$
$|\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda_{t}))|\leq\lambda_{t}$ $\quad~{}\text{if}~{}\widetilde{\beta}_{j}=0.$
@@
$=\lambda\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{X}_{\mathcal{A}}^{\top})^{+}\bm{\eta}_{\mathcal{A}}$
$=\lambda\bm{\eta}_{\mathcal{A}},$
where we used the identity $\mathbf{A}^{\top}\mathbf{A}\mathbf{A}^{+}=\mathbf{A}^{\top}$ and the last equality holds as $\bm{\eta}_{\mathcal{A}}\in\operatorname{row}(\mathbf{X}_{\mathcal{A}})$. Since
@@
$\lambda(1-\epsilon)\leq\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda))\operatorname{sign}(\widetilde{\beta}_{j}(\lambda))\leq\lambda$
for $j\in\mathcal{A}$ as required. Therefore, as $\lambda$ decreases, one of the following conditions must break: either $\|\mathbf{X}_{\mathcal{I}}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda))\|_{\infty}\leq\lambda$ or $\eta_{j}\neq 0$ for all $j\in\mathcal{A}$. The first breaks at the next joining time $\lambda^{\rm join}_{t+1}=\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}$, and second breaks at the next crossing time $\lambda_{t+1}^{\rm cross}=\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\}$. Since we only decrease $\lambda$ to $\lambda_{t+1}=\min\{\lambda_{t},\max\{{\lambda}_{t+1}^{\rm cross},\widetilde{\lambda}_{t+1}^{\rm join}\}\}$, we have verified that $\widetilde{\bm{\beta}}(\lambda)$ satisfies $\mathsf{KKT}_{\lambda}(\epsilon)$ for $\lambda\geq\lambda_{t+1}$.
@@
$=\mathbf{X}_{\mathcal{A}^{\ast}}^{+}\mathbf{X}_{\mathcal{A}^{\ast}}\widetilde{\bm{\beta}}_{\mathcal{A}^{\ast}}(\lambda_{t+1}),$
where we used the identity $\mathbf{A}^{+}(\mathbf{A}^{+})^{\top}\mathbf{A}^{\top}=\mathbf{A}^{+}$. Therefore, the solution $\widetilde{\bm{\beta}}_{\mathcal{A}^{\ast}}(\lambda_{t+1})$ to the above equation must be
@@
$\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}^{\ast}}\widetilde{\bm{\beta}}^{\ast}_{\mathcal{A}^{\ast}}(\lambda_{t+1}))=\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1}))$
for all $j\in[d]$, and so $\widetilde{\bm{\beta}}^{\ast}(\lambda_{t+1})$ satisfies the $\mathsf{KKT}_{\lambda_{t+1}}(\epsilon)$ condition.
@@
$=\mathbf{X}_{\mathcal{A}^{\ast}}^{+}\mathbf{X}_{\mathcal{A}^{\ast}}\widetilde{\bm{\beta}}_{\mathcal{A}^{\ast}}(\lambda_{t+1}),$
where we used the identity $\mathbf{A}^{+}(\mathbf{A}^{+})^{\top}\mathbf{A}^{\top}=\mathbf{A}^{+}$. Therefore, the solution $\widetilde{\bm{\beta}}_{\mathcal{A}^{\ast}}(\lambda_{t+1})$ to the above equation must be
@@
$\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}^{\ast}}\widetilde{\bm{\beta}}^{\ast}_{\mathcal{A}^{\ast}}(\lambda_{t+1}))=\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1}))$
@@
for all $j\in[d]$. It only remains to prove that the variable $\widetilde{i}_{t+1}^{\rm join}$ satisfies the $\mathsf{KKT}_{\lambda_{t+1}}(\epsilon)$ condition once it is added to $\mathcal{A}$. Indeed, first define $\Delta_{t+1}^{\rm join}\triangleq\min\{\lambda_{t},\widetilde{\lambda}^{\rm join}_{t+1}\}-\lambda_{t+1}^{\rm join}$ and notice that $\Delta_{t+1}^{\rm join}\leq\widetilde{\lambda}^{\rm join}_{t+1}\epsilon/2$ since $\lambda_{t}\geq\lambda_{t+1}^{\rm join}$ and $(1-\epsilon/2)\widetilde{\lambda}_{t+1}^{\rm join}\leq\lambda_{t+1}^{\rm join}$. Then, at the point $\lambda_{t+1}=\min\{\lambda_{t},\widetilde{\lambda}^{\rm join}_{t+1}\}$,
@@
$\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\widetilde{\lambda}^{\rm join}_{t+1}))\big{|}$ $=\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda^{\rm join}_{t+1}))+\Delta_{t+1}^{\rm join}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}}\big{|}$
$\geq\lambda_{t+1}^{\rm join}-\frac{\epsilon\widetilde{\lambda}^{\rm join}_{t+1}}{2}\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}}\big{|}$
@@
$=\lambda_{t+1}^{\rm join}-\frac{\epsilon\widetilde{\lambda}^{\rm join}_{t+1}}{2\lambda_{t}}\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\big{|}$
@@
$\geq\lambda_{t+1}^{\rm join}-\frac{\epsilon\widetilde{\lambda}^{\rm join}_{t+1}}{2\lambda_{t}}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}\|_{1}\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}$
@@
$\geq\lambda_{t+1}^{\rm join}-\frac{\epsilon\widetilde{\lambda}^{\rm join}_{t+1}}{2}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}\|_{1}$

### v2

@@ 24      $t\leftarrow t+1$
return coefficients $[(\lambda_{0},\widetilde{\bm{\beta}}(\lambda_{0})),(\lambda_{1},\widetilde{\bm{\beta}}(\lambda_{1})),\dots]$
@@
We now present our approximate quantum LARS Algorithm 3 for the pathwise Lasso. Its main idea is quite simple: we improve the search of the joining variable $i_{t+1}^{\rm join}$ and time $\lambda_{t+1}^{\rm join}$ over $i\in\mathcal{I}$ by approximately computing the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and using the approximate quantum minimum-finding subroutine from Chen and de Wolf [23] (Fact 14). More specifically, the joining variable $\widetilde{i}_{t+1}^{\rm join}$ is obtained to be any variable $j\in\mathcal{I}$ that maximises $\Lambda_{j}^{\rm join}(\mathcal{A})$ up to some small relative error $\epsilon$, i.e., at any iteration, Algorithm 3 randomly samples from the set
$\left\{j\in\mathcal{I}:\Lambda_{j}^{\rm join}(\mathcal{A})\geq\Big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\Big{)}\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\right\},$ (11)
where $\alpha_{\mathcal{A}}>0$ is such that $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\leq\alpha_{\mathcal{A}}$. Since the corresponding joining time $\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A})$ can be smaller than the true joining time $\lambda_{t+1}^{\rm join}=\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}$, we rescale it by a factor $\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}^{-1}$ and set $\widetilde{\lambda}_{t+1}^{\rm join}=\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}^{-1}\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A})$. This ensures that $\widetilde{\lambda}_{t+1}^{\rm join}\geq\lambda_{t+1}^{\rm join}$ and consequently that the current iteration stops before $\lambda_{t+1}^{\rm join}$, which guarantees that the approximate KKT conditions are satisfied as shown in Theorem 28 below.
We notice that, since the joining variable is randomly sampled from the set in Equation 11, it might be possible that $\widetilde{\lambda}_{t+2}^{\rm join}>\widetilde{\lambda}_{t+1}^{\rm join}$, i.e., the joining time at some later iteration is greater than the joining time at some earlier iteration. While this is counter to the original LARS algorithm, where the regularisation parameter $\lambda$ always decreases, there is in principle no problem in allowing $\lambda$ to increase within a small interval as long as the approximate KKT conditions are satisfied. Nonetheless, in order to avoid redundant segments in the Lasso path, we set the next iteration’s regularisation parameter $\lambda_{t+1}$ as the minimum between the current $\lambda_{t}$ and $\max\{\lambda_{t+1}^{\rm cross},\widetilde{\lambda}_{t+1}^{\rm join}\}$. The solution path can thus stay “stationary” for a few iterations. Ideally, one could just move all the variables from the set in Equation 11 into $\mathcal{A}$ at once and the result would be the same. This is reminiscent to the approximate homotopy algorithm of Mairal and Yu [78].
@@
Let $T\in\mathbb{N}$, $\epsilon\geq 0$, $\mathbf{X}\in\mathbb{R}^{n\times d}$, and $\mathbf{y}\in\mathbb{R}^{n}$. For $\mathcal{A}\subseteq[d]$ of size $|\mathcal{A}|\leq T$, let $\alpha_{\mathcal{A}}>0$ be such that $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\leq\alpha_{\mathcal{A}}$. Consider an approximate LARS algorithm (e.g., Algorithm 3) that returns a solution path $\widetilde{\mathcal{P}}=\{\widetilde{\bm{\beta}}(\lambda)\in\mathbb{R}^{d}:\lambda>0\}$ with at most $T$ kinks and wherein, at each iteration $t$, the joining variable $\widetilde{i}_{t+1}^{\rm join}$ is taken from the set
$\left\{j\in\mathcal{I}:\Lambda_{j}^{\rm join}(\mathcal{A})\geq\Big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\Big{)}\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\right\}$
and the corresponding joining time $\widetilde{\lambda}_{t+1}^{\rm join}$ is
$\widetilde{\lambda}_{t+1}^{\rm join}=\Big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\Big{)}^{-1}\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A}),$
@@
$\Lambda_{j}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{j}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}},\quad\bm{\mu}=\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y},\quad\bm{\theta}=(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}},\quad\bm{\eta}_{\mathcal{A}}=\frac{1}{\lambda_{t}}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t})).$
@@
The proof is by induction on the iteration loop $t$. The case $t=0$ is trivial as $\hat{\bm{\beta}}(\lambda)=\mathbf{0}$ for $\lambda\geq\lambda_{0}$. Assume then that the computed path through iteration $t-1$ satisfies $\mathsf{KKT}_{\lambda}(\epsilon)$ for all $\lambda\geq\lambda_{t}$. Consider the $t$-th iteration with active set $\mathcal{A}=\{i\in[d]:\widetilde{\beta}_{i}\neq 0\}$ and equicorrelation signs $\bm{\eta}_{\mathcal{A}}=\frac{1}{\lambda_{t}}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda_{t}))$. The induction hypothesis implies that the current Lasso solution $\widetilde{\bm{\beta}}(\lambda_{t})$ satisfies $\mathsf{KKT}_{\lambda_{t}}(\epsilon)$ at $\lambda_{t}$:
$\lambda_{t}(1-\epsilon)\leq\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda_{t}))\operatorname{sign}(\widetilde{\beta}_{j}(\lambda_{t}))\leq\lambda_{t}$ $\quad\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \widetilde{\beta}_{j}\neq 0,$
$|\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda_{t}))|\leq\lambda_{t}$ $\quad\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \widetilde{\beta}_{j}=0.$
@@
$=\lambda\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{X}_{\mathcal{A}}^{\top})^{+}\bm{\eta}_{\mathcal{A}}$ (by $\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}}\mathbf{X}^{+}_{\mathcal{A}}=\mathbf{X}_{\mathcal{A}}^{\top}$)
$=\lambda\bm{\eta}_{\mathcal{A}},$ (by $\bm{\eta}_{\mathcal{A}}\in\operatorname{row}(\mathbf{X}_{\mathcal{A}})$)
where we used the identity $\mathbf{A}^{\top}\mathbf{A}\mathbf{A}^{+}=\mathbf{A}^{\top}$. Since
@@
$\lambda(1-\epsilon)\leq\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda))\operatorname{sign}(\widetilde{\beta}_{j}(\lambda))\leq\lambda\quad\text{for all}\leavevmode\nobreak\ j\in\mathcal{A}$
as required. As $\lambda$ decreases, one of the following conditions must break: either $\eta_{j}\neq 0$ for all $j\in\mathcal{A}$ or $\|\mathbf{X}_{\mathcal{I}}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}(\lambda))\|_{\infty}\leq\lambda$. The first breaks at the next crossing time $\lambda_{t+1}^{\rm cross}=\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\}$, and the second breaks at the next joining time $\lambda^{\rm join}_{t+1}=\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}$. Since we only decrease $\lambda$ to $\lambda_{t+1}=\min\{\lambda_{t},\max\{{\lambda}_{t+1}^{\rm cross},\widetilde{\lambda}_{t+1}^{\rm join}\}\}$, we have verified that $\widetilde{\bm{\beta}}(\lambda)$ satisfies $\mathsf{KKT}_{\lambda}(\epsilon)$ for $\lambda\geq\lambda_{t+1}$.
@@
$=\mathbf{X}_{\mathcal{A}^{\ast}}^{+}\mathbf{X}_{\mathcal{A}^{\ast}}\widetilde{\bm{\beta}}_{\mathcal{A}^{\ast}}(\lambda_{t+1}),$ (by $\mathbf{X}_{\mathcal{A}^{\ast}}^{+}(\mathbf{X}_{\mathcal{A}^{\ast}}^{+})^{\top}\mathbf{X}_{\mathcal{A}^{\ast}}^{\top}=\mathbf{X}_{\mathcal{A}^{\ast}}^{+}$)
again using that $\mathbf{A}^{+}(\mathbf{A}^{+})^{\top}\mathbf{A}^{\top}=\mathbf{A}^{+}$. The solution $\widetilde{\bm{\beta}}_{\mathcal{A}^{\ast}}(\lambda_{t+1})$ to the above equation must be
@@
$\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}^{\ast}}\widetilde{\bm{\beta}}^{\ast}_{\mathcal{A}^{\ast}}(\lambda_{t+1}))=\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1}))\quad\text{for all}\leavevmode\nobreak\ j\in[d]$
and so $\widetilde{\bm{\beta}}^{\ast}(\lambda_{t+1})$ satisfies the $\mathsf{KKT}_{\lambda_{t+1}}(\epsilon)$ condition.
@@
$=\mathbf{X}_{\mathcal{A}^{\ast}}^{+}\mathbf{X}_{\mathcal{A}^{\ast}}\widetilde{\bm{\beta}}_{\mathcal{A}^{\ast}}(\lambda_{t+1}),$ (by $\mathbf{X}_{\mathcal{A}^{\ast}}^{+}(\mathbf{X}_{\mathcal{A}^{\ast}}^{+})^{\top}\mathbf{X}_{\mathcal{A}^{\ast}}^{\top}=\mathbf{X}_{\mathcal{A}^{\ast}}^{+}$)
again using that $\mathbf{A}^{+}(\mathbf{A}^{+})^{\top}\mathbf{A}^{\top}=\mathbf{A}^{+}$. The solution $\widetilde{\bm{\beta}}_{\mathcal{A}^{\ast}}(\lambda_{t+1})$ to the above equation must be
@@
$\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}^{\ast}}\widetilde{\bm{\beta}}^{\ast}_{\mathcal{A}^{\ast}}(\lambda_{t+1}))=\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1}))\quad\text{for all}\leavevmode\nobreak\ j\in[d].$
@@
It only remains to prove that the variable $\widetilde{i}_{t+1}^{\rm join}$ satisfies the $\mathsf{KKT}_{\lambda_{t+1}}(\epsilon)$ condition once it is added to $\mathcal{A}$. Indeed, first notice that the variable $\widetilde{i}_{t+1}^{\rm join}$ achieves correlation
$\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}))|=\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}$
at the point $\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}=\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A})$. Define $\Delta_{t+1}^{\rm join}\triangleq\min\{\lambda_{t},\widetilde{\lambda}^{\rm join}_{t+1}\}-\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}$ and note that $\Delta_{t+1}^{\rm join}\leq\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\widetilde{\lambda}^{\rm join}_{t+1}$ since $\lambda_{t}\geq\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}$ and $\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}\widetilde{\lambda}_{t+1}^{\rm join}\leq\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}$. Then, at the point $\lambda_{t+1}=\min\{\lambda_{t},\widetilde{\lambda}^{\rm join}_{t+1}\}$,
$\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\widetilde{\lambda}^{\rm join}_{t+1}))\big{|}$ $=\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}))+\Delta_{t+1}^{\rm join}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}}\big{|}$
@@
$\geq\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}))\big{|}-\Delta_{t+1}^{\rm join}\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}}\big{|}$ (Triangle inequality)
$=\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}-\Delta_{t+1}^{\rm join}\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}}\big{|}$ ($\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}))\big{|}=\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}$)
@@
$=\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}-\frac{\Delta_{t+1}^{\rm join}}{\lambda_{t}}\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\big{|}$
@@
$\geq\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}-\frac{\Delta_{t+1}^{\rm join}}{\lambda_{t}}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}\|_{1}\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}$ (Hölder’s inequality)
@@
$\geq\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}-\Delta_{t+1}^{\rm join}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}\|_{1}$ ($\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}\leq\lambda_{t}$)